# Gravity Is Induced By Renormalization Group Flow

### v1

@@ I Introduction
To this end, we use the holographic renormalization deBoer:1999tgo ; deHaro:2000vlm ; Bianchi:2001kw , i.e., the RG formulated within the AdS/CFT duality Maldacena:1997re , also known as holography Witten:1998qj ; Aharony:1999ti . AdS/CFT and its most commonly used limit, the gauge/gravity correspondence, provides a framework to formulate $d$-dimensional QFT in terms of a gravity theory on $(d+1)$-dimensional anti-de Sitter AdSd+1 spacetime. In this correspondence, the boundary QFT is encoded in a classical gravity theory in the bulk, with the radial (or holographic) direction in AdS playing the role of the RG scale Skenderis:2002wp . Radial evolution in the bulk then corresponds to RG flow in the boundary theory Sheikh-Jabbari:2013AdSCFT . While the arguments and results are by and large generic, here we focus on the $d=4$ case, a $4d$ QFT dual to Einstein gravity on AdS5.
@@ III Renormalization Group Induced Gravity
The above is the deformation flow equation for the boundary theory, with $r$ serving as the RG flow parameter. It is a first-order differential equation that admits a unique solution, once one provides the boundary (or initial) condition,
@@
The right-hand side of (14) denotes the boundary action evaluated at the asymptotic AdS5 boundary $\Sigma$ defined according to the standard gauge/gravity duality with Dirichlet bc. Eq. (13) captures the central statement that deforming the boundary theory by the TT operator enables the construction of finite-cutoff holography McGough:2016lol ; Taylor:2018xcy ; Hartman:2018tkw ; Parvizi:2025shq ; Parvizi:2025wsg .
@@
$\hskip-11.38109pt\begin{split}&{S}_{\text{bdy}}\Big{|}_{\Sigma(r)}={S}_{\text{bdy}}\Big{|}_{\Sigma_{0}}+\frac{1}{2\kappa_{4}(\mu)}\int_{\Sigma(r)}\sqrt{-\gamma}\Bigg{[}R[\gamma]-2\Lambda_{4}(\mu)\hskip-11.38109pt\\ &+2\kappa_{4}(\mu)\beta(\mu)\bigg{(}R_{ab}[\gamma]R^{ab}[\gamma]-\frac{1}{3}R[\gamma]^{2}\bigg{)}\Bigg{]}+\mathcal{O}(\mu^{-4}),\end{split}\hskip-2.84526pt$ (16)
with
[…]
where $\Sigma_{0}=\Sigma(r_{0})$, $\kappa_{4}(\mu)$ is the effective induced $4d$ Newton constant, $\Lambda_{4}(\mu)$ is the effective $4d$ cosmological constant and ${\gamma}_{ab}(x^{a};\mu)$ is the renormalized $4d$ gravitational metric. The $\beta$-term denotes the logarithmic corrections to the gravity action. Note that the effective $4d$ gravity couplings (17) have been obtained by assuming the following boundary conditions on the RGF equations: $\kappa_{4}=0,\Lambda=0,\beta=0$ at $\Sigma_{0}$ (large $\mu_{0}\gtrsim\mu$).
@@
While we arrived at the above result using a holographic analysis and through a $5d$ gravity description, the above could be completely discussed without using the $5d$ gravity and only within a $4d$ gauge theory and RG flow there: We start with a non-gravitating $4d$ gauge theory in the UV, flowing to the IR, among the irrelevant couplings turned on, there is TT, cf. (9) and (13), which is universal to all $4d$ QFTs and shows up as a $4d$ gravity theory. The above is already beyond the usual Wilsonian EFT picture, because the $4d$ metric is now viewed as an independent dynamical field. The $5d$ holographic description prompts the idea that there is no fundamental reason not to allow for the $4d$ metric to be turned on as an effective $4d$ dynamical field. The fact that this is not only a valid viewpoint, but indeed is required, can be understood within the holographic framework, as we discuss in the next section.
IV RG flow of boundary conditions
@@
$\delta h_{ab}(\epsilon)=-\frac{\kappa_{5}}{5L^{3}}\,\delta\left(L^{4}T_{ab}(\epsilon)\right)+\mathcal{O}(\epsilon^{2})\,$ (18)
@@ V Outlook
While we used $5d$ holographic setting for our computations and analysis, the statements and the results can be stated only within the $4d$ theory and its RG flow, without recrossing to $5d$.
@@
The RG-induced gravity arises from two features: (1) RG flow in boundary conditions, (18) and (20). As we fix the Dirichlet bc on the AdS boundary (and hence start with a non-gravitating theory in the UV), the bc for the metric at any radius inside AdS is a mixture of Neumann and Dirichlet, and hence the boundary metric away from the UV admits non-trivial fluctuations. (2) Induced Einstein gravity. Variation of the on-shell bulk action involves a TT term, cf. (13). Upon the gauge/gravity correspondence and recalling the $5d$ “Hamiltonian constraint” (8a), this may be replaced with an effective $4d$ Einstein gravity plus curvature squared corrections (16).
@@
Figure 2: RG flow in the boundary theory, starting from a UV CFT on a flat background. Along the flow, and for $\mu>\mu_{\text{\tiny{DT}}}$ with the transmutation scale $\mu_{\text{\tiny{DT}}}$, the theory admits two equivalent interpretations: a QFT deformed by the TT operator, or a gravitating boundary theory. For $\mu<\mu_{\text{\tiny{DT}}}$ the bc effectively becomes Neumann and only the gravitational interpretation remains valid.
@@
In the analysis of the main text, we considered an unrenormalized holographic RG flow setting. However, it is known that the on-shell gravity/boundary action is divergent. In the holographic renormalization program deBoer:1999tgo ; deHaro:2000vlm , appropriate counterterms have been introduced to remove the divergences Balasubramanian:1999re ; Emparan:1999pm ; Skenderis:2002wp . In the appendix D, we have extended our analysis to the renormalized case. One finds a “holographically renormalized induced $4d$ gravity” which, compared to the unrenormalized case, has the following remarkable features: (1) Effective (renormalized) $4d$ gravity coupling is now becoming a constant, with no scale $\mu$ dependence. (2) The effective (renormalized) cosmological constant is zero, prompting some ideas to resolve the cosmological constant problem.
We close with a comment that besides Einstein gravity, (16) has a logarithmic correction by the curvature-squared term, $R_{ab}R^{ab}-\frac{1}{3}R^{2}$. Notably, this combination corresponds to the critical gravity Lu:2011zk ; Kan:2013moa . There are other higher-order corrections to (16), involving third-order or higher powers or higher derivatives of curvature. There are also corrections involving products of second or higher powers of curvature and energy-momentum tensor of the matter fields coupled to the effective $4d$ gravity. Our holographic framework allows for a systematic derivation of these corrections. See appendices C and D for a sketch of the derivation.
@@ Acknowledgement
We would like to thank Daniel Grumiller, Juan Maldacena, and Steve Shenker for comments on the manuscript. MMShJ acknowledges BIMSA and Hossein Yavartanoo for the hospitality, where he gave a minicourse on the Freelance Holography Program, and Sergio Cecotti for a comment in the minicourse that sparked this work. The work of HA is supported by the Beijing Natural Science Foundation under Grant No. IS23018. The work of VT is supported by the Iran National Science Foundation (INSF) under project No. 4040771.
@@
In the appendices A, B and C, we provide details of derivations. In the main text we discussed the gravity induced on the boundary within an unregularized holographic RG flow analysis. However, one could have used the regularized holographic RG flow Skenderis:2002wp ; Balasubramanian:1999re ; Emparan:1999pm .
@@ Appendix A Details of some derivations
where ${\cal G}^{abcd}$ is the WdW metric,
@@ Appendix C Holographically induced gravity action
To solve the above differential equation, consider the following solution anstaz:
@@
and solve for the unknown coefficient $\kappa_{4}(r),\Lambda_{4}(r),\beta(r)$. Note that in the above ${S}_{\text{bdy}}\Big{|}_{\Sigma_{0}}$is an integration constant, for a boundary at a large $r=r_{0}$, and we solve the equation for $r<r_{0}$. To solve this equation, we need to have the information of $\gamma_{ab}(r)$. Recalling the general form of $5d$ Einstein equations, we have:
@@
where $S_{ab}[\gamma]$} is the dimensionless Schouten tensor. The above is the leading part of the exact equation (46aw). In the $4d$ theory, the above is to be viewed as the wave-function renormalization for the metric field $\gamma_{ab}(x^{a};r)$. In a different viewpoint, (40) may be viewed as a Ricci (Schouten) flow equation.
@@ D.1 Renormalized action principle
$\mathcal{T}\hskip-2.84526pt{\mathcal{T}}=\mathcal{T}^{ab}\mathcal{T}_{ab}-\frac{1}{3}\mathcal{T}^{2}\,.$ (46au)
@@
$\begin{split}r\partial_{r}\mathcal{T}&=-2\mathcal{T}+L\kappa_{5}\Big{(}3\mathcal{T}\hskip-2.84526pt{\mathcal{T}}+\frac{\mathcal{T}^{2}}{3}\Big{)}+(4S^{ab}\mathcal{T}_{ab}-S\mathcal{T})\\ &+\frac{1}{L\kappa_{5}}(S_{ab}S^{ab}-S^{2})\,.\end{split}$ (46av)
@@ D.2 RG equation for the boundary metric $\gamma_{ab}(x^{c};r)$
$\hskip-5.69046ptr\partial_{r}\gamma_{ab}=\frac{2L^{2}}{r^{2}}S_{ab}[\gamma]+\frac{2L^{5}\kappa_{5}}{r^{4}}\left(\hat{\mathcal{T}}_{ab}-\frac{\hat{\mathcal{T}}}{3}\gamma_{ab}\right)\,.$ (46aw)
@@
$\hat{\mathcal{T}}_{ab}:=\left(\frac{r}{L}\right)^{2}\mathcal{T}_{ab}\,.$ (46ax)
In the large-$r$ limit, the second term becomes subleading, and the dominant contribution to the RG flow of
$r\partial_{r}\gamma_{ab}=\frac{2L^{2}}{r^{2}}S_{ab}[\gamma]+{\cal O}\left(\left(\frac{L}{r}\right)^{4}\right)\,.$ (46ay)
This equation closely resembles the celebrated Ricci flow equation, suggesting a geometric interpretation of the boundary RG flow in terms of a curvature-driven evolution. This is (40), which we used in Appendix C.
@@ D.3 Renormalized deformation flow equation
${S}^{\text{\tiny{Ren}}}_{r}[g]\Big{|}_{\text{on-shell}}={S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}\,$, (46az)
@@
$\lim_{r\to\infty}{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}={S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma}\,$, (46ba)
@@
$\begin{split}r\frac{\mathrm{d}}{\mathrm{d}r}{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}&=\int_{\Sigma(r)}\sqrt{-h}\Big{[}\mathcal{T}+S^{ab}\mathcal{T}_{ab}+\frac{1}{L\kappa_{5}}(S^{ab}S_{ab}-S^{2})\Big{]}\\ &=\int_{\Sigma(r)}\sqrt{-\gamma}\Big{[}\hat{\mathcal{T}}+\frac{L^{2}}{r^{2}}S^{ab}[\gamma]\hat{\mathcal{T}}_{ab}\\ &+\frac{1}{L\kappa_{5}}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}\Big{]}.\end{split}$ (46bb)
@@
$\begin{split}&\kappa_{5}\,\frac{L^{4}}{r^{4}}\hat{\mathcal{T}\hskip-2.84526pt{\mathcal{T}}}+2\frac{\hat{\mathcal{T}}}{L}+2\frac{L}{r^{2}}S^{ab}[\gamma]\hat{\mathcal{T}}_{ab}\\ &+\frac{1}{L^{2}\kappa_{5}}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}=0\,.\end{split}$ (46bc)
@@
$\hat{\mathcal{T}}=-\frac{1}{2L\kappa_{5}}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}\quad\text{on $\Sigma$}\,.$ (46bd)
@@
$\begin{split}\hskip-5.69046ptr\frac{\mathrm{d}}{\mathrm{d}r}{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}=&\frac{1}{2L\kappa_{5}}\int_{\Sigma(r)}\sqrt{-\gamma}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}\\ &+\mathcal{O}(r^{-4})\,.\end{split}$ (46be)
To solve this equation, we again use the solution ansatz (39), then we find the following RG equations
$\begin{split}&\frac{\operatorname{d}\!{}}{\operatorname{d}\!{}r}\left(\frac{\Lambda_{4}}{\kappa_{4}}\right)=0\,,\quad\frac{\operatorname{d}\!{}}{\operatorname{d}\!{}r}\left(\frac{1}{\kappa_{4}}\right)=\frac{L}{3}\left(\frac{L}{r}\right)^{3}\frac{\Lambda_{4}}{\kappa_{4}}\,,\\ &\frac{\operatorname{d}\!{}}{\operatorname{d}\!{}r}\beta=\frac{L^{3}}{8\kappa_{5}r}+\frac{L}{2}\left(\frac{L}{r}\right)^{3}\frac{1}{\kappa_{4}}\,.\end{split}$ (46bf)
@@
$\begin{split}&{\Lambda_{4}}=0\,,\quad\kappa_{4}=\mathrm{const}\,,\\ &\beta=\frac{L^{3}}{8\kappa_{5}}\ln\left(\frac{r}{r_{0}}\right)-\frac{L^{2}}{4\kappa_{4}}\left(\frac{L}{r}\right)^{2}\,.\end{split}$ (46bg)
@@
$\begin{split}&{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}={S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma_{0}}+\frac{1}{2\kappa_{4}}\int_{\Sigma(r)}\sqrt{-\gamma}R[\gamma]\\ &+\beta(r)\int_{\Sigma(r)}\sqrt{-\gamma}\big{(}R_{ab}[\gamma]R^{ab}[\gamma]-\frac{1}{3}R^{2}[\gamma]\big{)}+\mathcal{O}(r^{-4}).\end{split}$ (46bh)
@@
    While in both cases the leading part of the coefficient of curvature-squared terms has the logarithmic scale dependence, in the renormalized case, there is a subleading piece that falls off quadratically with scale.

### v2

@@ I Introduction
To this end, we use the holographic renormalization deBoer:1999tgo ; deHaro:2000vlm ; Bianchi:2001kw ; Akhmedov:2010mz , i.e., the RG formulated within the AdS/CFT duality Maldacena:1997re , also known as holography Witten:1998qj ; Aharony:1999ti . AdS/CFT and its most commonly used limit, the gauge/gravity correspondence, provides a framework to formulate $d$-dimensional QFT in terms of a gravity theory on $(d+1)$-dimensional anti-de Sitter AdSd+1 spacetime. In this correspondence, the boundary QFT is encoded in a classical gravity theory in the bulk, with the radial (or holographic) direction in AdS playing the role of the RG scale Skenderis:2002wp . Radial evolution in the bulk then corresponds to RGF in the boundary theory Sheikh-Jabbari:2013AdSCFT . While the arguments and results are by and large generic, here we focus on the $d=4$ case, a $4d$ QFT dual to Einstein gravity on AdS5.
@@ III Renormalization Group Induced Gravity
The above is the deformation flow equation for the boundary theory, with $r$ serving as the RGF parameter. It is a first-order differential equation that admits a unique solution, once one provides the boundary (or initial) condition,
@@
The right-hand side of (14) denotes the boundary action evaluated at the asymptotic AdS5 boundary $\Sigma$ defined according to the standard gauge/gravity duality with Dirichlet bc. Eq. (13) captures the central statement that deforming the boundary theory by the TT operator enables the construction of finite-cutoff holography McGough:2016lol ; Taylor:2018xcy ; Hartman:2018tkw ; Parvizi:2025shq ; Parvizi:2025wsg ; Babaei-Aghbolagh:2024hti ; He:2025ppz .
@@
$\hskip-11.38109pt\begin{split}{S}_{\text{bdy}}\Big{|}_{\Sigma(r)}=&{S}_{\text{bdy}}\Big{|}_{\Sigma_{0}}+\frac{1}{2\kappa_{4}(\mu)}\int_{\Sigma(r)}\sqrt{-\gamma}\Bigg{[}R[\gamma]-2\Lambda_{4}(\mu)\hskip-11.38109pt\\ &+{\kappa_{4}(\mu)\beta(\mu)W_{abcd}[\gamma]W^{abcd}[\gamma]}\Bigg{]}+\mathcal{O}(\mu^{-4}),\end{split}\hskip-5.69054pt$ (16)
where $W_{abcd}[\gamma]$ is the Weyl curvature of metric $\gamma_{ab}$ and
$\begin{split}\mu:=\frac{r}{L^{2}},\qquad&\kappa_{4}(\mu)=\frac{2\kappa_{5}}{L^{3}}\frac{1}{\mu^{2}}\,,\\ \Lambda_{4}(\mu):=6\mu^{2}\left(\frac{\mu_{0}^{4}}{\mu^{4}}-1\right),\quad&\beta(\mu)=\frac{2\kappa_{5}}{4L^{3}}\ln\left(\frac{\mu}{\mu_{0}}\right)\,.\end{split}$ (17)
Here $\Sigma_{0}=\Sigma(r_{0})$, $\kappa_{4}(\mu)$ is the effective induced $4d$ Newton constant, $\Lambda_{4}(\mu)$ is the effective $4d$ cosmological constant and ${\gamma}_{ab}(x^{a};\mu)$ is the renormalized $4d$ gravitational metric. The $\beta$-term denotes the logarithmic corrections to the Einstein gravity by the seminal conformal gravity Stelle:1976gc ; Adler:1982ri ; Mannheim:2011ds ; tHooft:2011aa ; Maldacena:2011mk . Note that the effective $4d$ gravity couplings (17) have been obtained by assuming the following boundary conditions on the RGF equations: $\kappa_{4}=0,\Lambda=0,\beta=0$ at $\Sigma_{0}$ (large $\mu_{0}\gtrsim\mu$).
@@
While we arrived at the above result using a holographic analysis and through a $5d$ gravity description, the above could be completely discussed without using the $5d$ gravity and only within a $4d$ gauge theory and RGF there: We start with a non-gravitating $4d$ gauge theory in the UV, flowing to the IR, among the irrelevant couplings turned on, there is TT, cf. (9) and (13), which is universal to all $4d$ QFTs and shows up as a $4d$ gravity theory. The above is already beyond the usual Wilsonian EFT picture, because the $4d$ metric is now viewed as an independent dynamical field. The $5d$ holographic description prompts the idea that there is no fundamental reason not to allow for the $4d$ metric to be turned on as an effective $4d$ dynamical field. The fact that this is not only a valid viewpoint, but indeed is required, can be understood within the holographic framework, as we discuss in the next section.
IV Renormalization flow of boundary conditions
@@
$\delta h_{ab}(\epsilon)=-\frac{\kappa_{5}}{5L^{3}}\,\delta\left(L^{4}T_{ab}(\epsilon)\right)+\mathcal{O}(\epsilon^{2})\,$, (18)
@@ V Outlook
While we used $5d$ holographic setting for our computations and analysis, the statements and the results can be stated only within the $4d$ theory and its RGF, without recrossing to $5d$.
@@
The RG-induced gravity arises from two features: (1) RGF in boundary conditions, (18) and (20). As we fix the Dirichlet bc on the AdS boundary (and hence start with a non-gravitating theory in the UV), the bc for the metric at any radius inside AdS is a mixture of Neumann and Dirichlet, and hence the boundary metric away from the UV admits non-trivial fluctuations. (2) Induced Einstein gravity. Variation of the on-shell bulk action involves a TT term, cf. (13). Upon the gauge/gravity correspondence and recalling the $5d$ “Hamiltonian constraint” (8a), this may be replaced with an effective $4d$ Einstein gravity plus curvature squared corrections (16).
@@
Figure 2: RGF in the boundary theory, starting from a UV CFT on a flat background. Along the flow, and for $\mu>\mu_{\text{\tiny{DT}}}$ with the transmutation scale $\mu_{\text{\tiny{DT}}}$, the theory admits two equivalent interpretations: a QFT deformed by the TT operator, or a gravitating boundary theory. For $\mu<\mu_{\text{\tiny{DT}}}$ the bc effectively becomes Neumann and only the gravitational interpretation remains valid.
@@
In the analysis of the main text, we considered an unrenormalized holographic RGF setting. However, it is known that the on-shell gravity/boundary action is divergent. In the holographic renormalization program deBoer:1999tgo ; deHaro:2000vlm , appropriate counterterms have been introduced to remove the divergences Balasubramanian:1999re ; Emparan:1999pm ; Skenderis:2002wp . In the appendix D, we have extended our analysis to the renormalized case. One finds a “holographically renormalized induced $4d$ gravity” which, compared to the unrenormalized case, has the following remarkable features: (1) Effective (renormalized) $4d$ gravity coupling is now becoming a constant, with no scale $\mu$ dependence. (2) The effective (renormalized) cosmological constant is zero, prompting some ideas to resolve the cosmological constant problem.
We close with a comment that besides Einstein gravity, (16) has a logarithmic correction by the Weyl-squared (conformal gravity) term Stelle:1976gc ; Adler:1982ri ; Mannheim:2011ds ; tHooft:2011aa ; Maldacena:2011mk . There are other higher-order corrections to (16), involving third-order or higher powers or higher derivatives of curvature. There are also corrections involving products of second or higher powers of curvature and energy-momentum tensor of the matter fields coupled to the effective $4d$ gravity. Our holographic framework allows for a systematic derivation of these corrections. See appendices C and D for a sketch of the derivation.
@@ Acknowledgement
We would like to thank Dionysios Anninos, Daniel Grumiller, Juan Maldacena, and Steve Shenker for comments on the manuscript. MMShJ acknowledges BIMSA and Hossein Yavartanoo for the hospitality, where he gave a minicourse on the Freelance Holography Program, and Sergio Cecotti for a comment in the minicourse that sparked this work. The work of HA is supported by the Beijing Natural Science Foundation under Grant No. IS23018. The work of VT is supported by the Iran National Science Foundation (INSF) under project No. 4040771.
@@
In the appendices A, B, and C, we provide details of derivations. In the main text we discussed the gravity induced on the boundary within an unregularized holographic RGF analysis. However, one could have used the regularized holographic RGF Skenderis:2002wp ; Balasubramanian:1999re ; Emparan:1999pm .
@@ Appendix A Details of some derivations
where ${\cal G}^{abcd}$ is the WdW metric PhysRev.160.1113 ,
@@ Appendix C Holographically induced gravity action
To solve the above differential equation, consider the following solution ansatz:
@@
and solve for the unknown coefficient $\kappa_{4}(r),\Lambda_{4}(r),\beta(r)$. Note that in writing the above, we used the identity,
[…]
where GB is the Gauss-Bonnet term which is a topological (total derivative) term in $4d$ and thus do not contribute to equations of motion. In the above ${S}_{\text{bdy}}\Big{|}_{\Sigma_{0}}$is an integration constant, for a boundary at a large $r=r_{0}$, and we solve the equation for $r<r_{0}$. To solve this equation, we need to have the information of $\gamma_{ab}(r)$. Recalling the general form of $5d$ Einstein equations, we have:
@@
where $S_{ab}[\gamma]$} is the dimensionless Schouten tensor. The above is the leading part of the exact equation (47ax). In the $4d$ theory, the above is to be viewed as the wave-function renormalization for the metric field $\gamma_{ab}(x^{a};r)$. In a different viewpoint, (41) may be viewed as a Ricci (Schouten) flow equation.
@@ D.1 Renormalized action principle
$\mathcal{T}\hskip-2.84526pt{\mathcal{T}}=\mathcal{T}^{ab}\mathcal{T}_{ab}-\frac{1}{3}\mathcal{T}^{2}\,.$ (47av)
@@
$\begin{split}r\partial_{r}\mathcal{T}&=-2\mathcal{T}+L\kappa_{5}\Big{(}3\mathcal{T}\hskip-2.84526pt{\mathcal{T}}+\frac{\mathcal{T}^{2}}{3}\Big{)}+(4S^{ab}\mathcal{T}_{ab}-S\mathcal{T})\\ &+\frac{1}{L\kappa_{5}}(S_{ab}S^{ab}-S^{2})\,.\end{split}$ (47aw)
@@ D.2 RG equation for the boundary metric $\gamma_{ab}(x^{c};r)$
$\hskip-5.69046ptr\partial_{r}\gamma_{ab}=\frac{2L^{2}}{r^{2}}S_{ab}[\gamma]+\frac{2L^{5}\kappa_{5}}{r^{4}}\left(\hat{\mathcal{T}}_{ab}-\frac{\hat{\mathcal{T}}}{3}\gamma_{ab}\right)\,.$ (47ax)
@@
$\hat{\mathcal{T}}_{ab}:=\left(\frac{r}{L}\right)^{2}\mathcal{T}_{ab}\,.$ (47ay)
In the large-$r$ limit, the second term becomes subleading, and the dominant contribution to the RGF of
$r\partial_{r}\gamma_{ab}=\frac{2L^{2}}{r^{2}}S_{ab}[\gamma]+{\cal O}\left(\left(\frac{L}{r}\right)^{4}\right)\,.$ (47az)
This equation closely resembles the celebrated Ricci flow equation, suggesting a geometric interpretation of the boundary RGF in terms of a curvature-driven evolution. This is (41), which we used in Appendix C.
@@ D.3 Renormalized deformation flow equation
${S}^{\text{\tiny{Ren}}}_{r}[g]\Big{|}_{\text{on-shell}}={S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}\,$, (47ba)
@@
$\lim_{r\to\infty}{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}={S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma}\,$, (47bb)
@@
$\begin{split}r\frac{\mathrm{d}}{\mathrm{d}r}{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}&=\int_{\Sigma(r)}\sqrt{-h}\Big{[}\mathcal{T}+S^{ab}\mathcal{T}_{ab}+\frac{1}{L\kappa_{5}}(S^{ab}S_{ab}-S^{2})\Big{]}\\ &=\int_{\Sigma(r)}\sqrt{-\gamma}\Big{[}\hat{\mathcal{T}}+\frac{L^{2}}{r^{2}}S^{ab}[\gamma]\hat{\mathcal{T}}_{ab}\\ &+\frac{1}{L\kappa_{5}}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}\Big{]}.\end{split}$ (47bc)
@@
$\begin{split}&\kappa_{5}\,\frac{L^{4}}{r^{4}}\hat{\mathcal{T}\hskip-2.84526pt{\mathcal{T}}}+2\frac{\hat{\mathcal{T}}}{L}+2\frac{L}{r^{2}}S^{ab}[\gamma]\hat{\mathcal{T}}_{ab}\\ &+\frac{1}{L^{2}\kappa_{5}}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}=0\,.\end{split}$ (47bd)
@@
$\hat{\mathcal{T}}=-\frac{1}{2L\kappa_{5}}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}\quad\text{on $\Sigma$}\,.$ (47be)
@@
$\begin{split}\hskip-5.69046ptr\frac{\mathrm{d}}{\mathrm{d}r}{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}=&\frac{1}{2L\kappa_{5}}\int_{\Sigma(r)}\sqrt{-\gamma}\big{(}S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}\big{)}\\ &+\mathcal{O}(r^{-4})\,.\end{split}$ (47bf)
To solve this equation, we recall $S^{ab}[\gamma]S_{ab}[\gamma]-S[\gamma]^{2}=\frac{L^{4}}{4}(R_{ab}[\gamma]R^{ab}[\gamma]-\frac{1}{3}R^{2}[\gamma])$ and (40), and again use the solution ansatz (39). We find the following RG equations
$\begin{split}&\frac{\operatorname{d}\!{}}{\operatorname{d}\!{}r}\left(\frac{\Lambda_{4}}{\kappa_{4}}\right)=0\,,\quad\frac{\operatorname{d}\!{}}{\operatorname{d}\!{}r}\left(\frac{1}{\kappa_{4}}\right)=\frac{L}{3}\left(\frac{L}{r}\right)^{3}\frac{\Lambda_{4}}{\kappa_{4}}\,,\\ &\frac{\operatorname{d}\!{}}{\operatorname{d}\!{}r}\beta=\frac{L^{3}}{8\kappa_{5}r}+\frac{L}{2}\left(\frac{L}{r}\right)^{3}\frac{1}{\kappa_{4}}\,.\end{split}$ (47bg)
@@
$\begin{split}&{\Lambda_{4}}=0\,,\quad\kappa_{4}=\mathrm{const}\,,\\ &\beta=\frac{L^{3}}{8\kappa_{5}}\ln\left(\frac{r}{r_{0}}\right)-\frac{L^{2}}{4\kappa_{4}}\left(\frac{L}{r}\right)^{2}\,.\end{split}$ (47bh)
@@
$\begin{split}&{S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma(r)}={S}^{\text{\tiny{Ren}}}_{\text{bdy}}\Big{|}_{\Sigma_{0}}+\frac{1}{2\kappa_{4}}\int_{\Sigma(r)}\sqrt{-\gamma}R[\gamma]\\ &+\frac{1}{2}\beta(r)\int_{\Sigma(r)}\sqrt{-\gamma}\ W^{abcd}[\gamma]W_{abcd}[\gamma]+\mathcal{O}(r^{-4})\,.\end{split}$ (47bi)
@@
    While in both cases the leading part of the coefficient of Weyl-squared (conformal gravity) terms has the logarithmic scale dependence, in the renormalized case, there is a subleading piece that falls off quadratically with scale.